# A Million Three-body Binaries Caught by Gaia

### v2

@@ B.2 The Cluster Initial Mass Function
First, let’s assert the form of the CIMF distribution
$f_{\rm CIMF}(M_{\rm cl0},t,R_{G})\propto M_{\rm cl0}^{-2}e^{-M_{\rm cl0}/M_{c}(t,R_{G})}$ (B4)
where $M_{c}(t,R_{G})$ is the time- and Galactocentric radius-dependent cutoff of the Schechter function. Considering that L. C. Johnson et al. (2017) empirically found that $M_{c}$ may be related to the star formation rate surface density, $\Sigma_{\rm SFR}$, with the expression
$\log_{10}\left(\frac{M_{c}}{M_{\odot}}\right)=1.07\log_{10}\left(\frac{\Sigma_{\rm SFR}}{M_{\odot}\rm{yr}^{-1}\rm{kpc}^{-2}}\right),$ (B5)
@@
$\Sigma_{\rm SFR}(t,R_{G})$ $=\frac{M_{d}+M_{b}}{2\pi R_{G}}f_{\rm ha}(R_{G})SFH_{\rm ha}(t)+\frac{M_{d}}{2\pi R_{G}}f_{\rm la}(R_{G})SFH_{\rm la}(t,R_{G})$ (B6)
which is related to $M_{c}(t,R_{G})$ using Equation (B5). The resulting normalization for Equation (B4) is
$f_{\rm CIMF}^{\rm norm}(\tilde{M}_{\rm norm})=\frac{\tilde{M}_{\rm norm}}{e^{-\tilde{M}_{\rm norm}}+\tilde{M}_{\rm norm}(0.1485+E_{1}(\tilde{M}_{\rm norm}))},\ \ \tilde{M}_{\rm norm}=M_{\rm min}/M_{c}(t,R_{G})$ (B7)
where $M_{\rm min}=20\,M_{\odot}$ is the smallest cluster mass we consider in the CIMF. Equation (B4) is normalized between $M_{\rm min}$ and $M_{c}(t,R_{G})$ since we consider $M_{c}$ to be the largest possible cluster mass in a star forming region (see also N. Choksi & J. M. D. Kruijssen, 2021).
@@ C.1.1 Wide Binary Disruption in the Disk
Binary disruption in the disk due to stellar encounters is not included a priori through phase mixing. Appending a disruption survival probability, $P_{\rm d,MW}$ (Equation C2), to our final equation (Equation D1) is an easy fix. The details of our $P_{\rm d,MW}$ derivation is discussed in § C.1.1. The resulting distribution in Equation (D1) returns the total number of binaries when integrated, but with initially unevolved binary samples. Their orbital elements remain identical to what was assigned at formation in the cluster; it is to this final distribution of surviving binaries we apply phase mixing, as described in § C.1.2.101010Note that phase mixing only modifies the eccentricity distribution of a binary population, leaving the SMA distribution unmodified.
The fraction of binaries which “diffuse” to a present-day Galactocentric orbit, $R_{G}^{\prime}$, will naturally be curtailed by strong scattering encounters in the disk. It is naturally preferable to consider an evolving disk model, accounting for the changing number and density of bodies in the disk with look-back time, $t^{\prime}$ (i.e., the time the binary is born and escapes the cluster). To avoid numerous intractable difficulties in the calculation, we instead adopt the conservative simplifying assumption of a constant disk density equal to the present-day density distribution.
@@
[…]
$v_{d}=\frac{R_{G}^{\prime}-R_{G}}{t^{\prime}},\ t^{\prime}=\tau_{\rm MW}-\tau-t$
where $v_{d}$ is the average “diffusion velocity” and $\tau$ is the age of the cluster when the binary is escapes.
@@
[…]
@@ C.1.2 Phase Mixing
After a binary escapes a cluster, if the binary is sufficiently small not to be destroyed by Galactic tides (i.e. if $a\lesssim 2\times 10^{5}$ AU), the binary evolves secularly under their influence (J. Heisler & S. Tremaine, 1986; C. Hamilton & R. R. Rafikov, 2019b, a; C. Hamilton, 2022; E. Grishin & H. B. Perets, 2022).
C. Hamilton (2022) showed that the consequence of such evolution is a “phase mixing” (PM) in the appropriate phase-space region. This secular evolution is described by the Galactic-tide Hamiltonian $\mathscr{H}$, given by (C. Hamilton & R. R. Rafikov, 2019b, a)
$\mathscr{H}$ $=\frac{\kappa^{2}a^{2}\mu}{8}\Big{[}\left(2+3\epsilon^{2}\right)\left(1-3\Gamma\cos^{2}i\right)-15\Gamma\epsilon^{2}\sin^{2}i\cos 2\omega\Big{]}$ (C4)
@@ C.2 Cumulative Scatter
where $(a^{\prime},\epsilon^{\prime 2})$ are the updated binary properties during time step $\delta t$, $m_{b}=1M_{\odot}$ is the total mass of the binary, $m_{p}=0.5M_{\odot}$ is the average stellar mass, $v_{d}$ is the average radial diffusion velocity of the binary—see also Equation (C2)—and $(\xi_{a},\xi_{\epsilon})$ are independent Gaussian random numbers with mean 0 and variance 1.111111Note that in C. Hamilton & S. Modak (2024), $\epsilon$ is the square of the binary eccentricity, while this work employs $\epsilon$ as the eccentricity due to the symbolic degeneracy with the Euler’s number.
@@
where the expression has been simplified to the case of $1M_{\odot}$ binaries with a relative velocity dispersion between GMC and disk stars of $\sim 30\rm{km\,s^{-1}}$. The probability that a binary evolving under CS survives being destroyed due to interactions with GMCs in each time-step is
@@
$M_{\rm cl0}$ $20\ M_{\odot}$ $10^{7}\ M_{\odot}$ an initial mass containing 40 bodies to a mass encompassing ${\gg}99\%$ of the parameter space
@@
$R_{G}^{\prime}$ $7.8\ \rm kpc$ $8.2\ \rm kpc$ the subset of final orbital radii within 200 pc of the sun
@@ Appendix D The Galactic 3BBF Rate
Having assembled the creation, diffusion, and destruction portions of the final distribution, we are finally in a position to discuss subtleties of the total integration. The final equation, our “Galactic 3BBF Rate,” (G3R) is the product of Equations A10, A12, A21, B3, B4, and C2, built with the numerous intermittent expressions. Written in functional form, the G3R is
$\mathcal{G}_{\rm MW}d\Omega=\frac{A_{*}}{A_{\rm R_{G,s}}}f_{\rm CIMF}(M_{\rm cl0},t,R_{G})\mathcal{C}(M_{\rm cl0},\tilde{\tau},R_{G},r,k,\epsilon,a)P_{\rm ie}(M_{\rm cl0},\tilde{\tau},R_{G},r,a)$ (D1)
×$\qquad\qquad\times\mathcal{DHL}(t,\tau,R_{G},R_{G}^{\prime})P_{\rm d,MW}(a,t,\tau,R_{G},R_{G}^{\prime})d\Omega,$
@@
where $A_{*}/A_{\rm R_{G,s}}=0.00625$ is the fraction of a ±$\pm 200\rm{pc}$ disk annulus at $R_{G,s}=8\rm{kpc}$ within a 200 pc radius of the Sun, and the substitutions $r=r_{t}\,\tilde{r}$, $dr=r_{t}d\tilde{r}$, $\tau=\tau_{\rm cl}\,\tilde{\tau}$, and $d\tau=\tau_{\rm cl}d\tilde{\tau}$ are made for a more convenient integration experience.
@@
The integration bounds are detailed in Table 3. A number of additional constraints are applied to Equation (D1), setting it to 00 when triggered. These constraints include:
@@
    $N_{\rm cl}(\tilde{\tau})\leq 10$; the validity of a Plummer model is contingent to many bodies being present. We arbitrarily select $N=10$ as a break point for integration.
@@
    $a\left(1-\epsilon\right)<0.1\rm{AU}$; collisional effects both before and after 3BBF would prevent such tight passages.
@@
    $\tau_{\rm MW}-\tau-t<0$; this only occurs if the migration timescale is larger than the time a binary has to migrate to $R_{G}^{\prime}$ from $R_{G}$ by the present day.
Incorporating all stated prior conditions with Equation (D1), the G3R is integrated over the entire parameter space using a custom written Monte Carlo integration scheme which divides the parameter space into $\sim 100$ sub-volumes and draws a total of $10^{9}$ samples, resolving the integral to high-precision. We additionally employ emcee (D. Foreman-Mackey et al., 2013) to uncover the underlying nine-dimensional distribution function accompanied by the same restrictions quoted above.121212emcee is run with 512 walkers (two per dimension) and is terminated when the fractional difference in autocorrelation time is $<1\%$ every $10^{4}$ steps. The resulting HDF files have a ${\sim}10\,\rm{GB}$ footprint.

### v3

@@ B.2 The Cluster Initial Mass Function
First, let’s assert the form of the CIMF distribution as found by A. Just et al. (2023)
$f_{\rm CIMF}(M_{\rm cl0},t,R_{G})\propto M_{\rm cl0}^{-1}(1+M_{\rm cl0}^{2.4})^{-1/2}e^{-M_{\rm cl0}/M_{c}(t,R_{G})}$ (B4)
where $M_{c}(t,R_{G})$ is the time- and Galactocentric radius-dependent cutoff of the Schechter function. L. C. Johnson et al. (2017) empirically found that $M_{c}$ may be related to the star formation rate surface density, $\Sigma_{\rm SFR}$, with the expression
$\log_{10}\left(\frac{M_{c}}{M_{\odot}}\right)=1.07\log_{10}\left(\frac{\Sigma_{\rm SFR}}{M_{\odot}\rm{yr}^{-1}\rm{kpc}^{-2}}\right).$ (B5)
@@
$\Sigma_{\rm SFR}(t,R_{G})$ $=\frac{M_{d}}{2\pi R_{G}}\left(f_{\rm ha}(R_{G})SFH_{\rm ha}(t)+f_{\rm la}(R_{G})SFH_{\rm la}(t,R_{G})\right),$ (B6)
which is related to $M_{c}(t,R_{G})$ using Equation (B5). The resulting normalization for Equation (B4) is numerically evaluated between $2M_{\odot}$ (A. Just et al., 2023) and $M_{c}(t,R_{G})$, since we consider $M_{c}$ to be the largest possible cluster mass in a star forming region (see also N. Choksi & J. M. D. Kruijssen, 2021).
@@ C.1.1 Wide Binary Disruption in the Disk
Binary disruption in the disk due to stellar encounters is not included a priori through phase mixing. Appending a disruption survival probability, $P_{\rm d,MW}$ (Equation C2), to our final equation is an easy fix. The final equation, Equation D1, returns the total number of binaries when integrated, but with initially unevolved binary samples. Their orbital elements remain identical to what was assigned at formation in the cluster—it is to this final distribution of surviving binaries we apply phase mixing, as described in § C.1.2.999Note that phase mixing only modifies the eccentricity distribution of a binary population, leaving the SMA distribution unmodified.
The fraction of binaries which “diffuse” to a present-day Galactocentric orbit, $R_{G}^{\prime}$, will naturally be curtailed by strong scattering encounters in the disk. It is naturally preferable to consider an evolving disk model, accounting for the changing number and density of bodies in the disk with look-back time, $t^{\prime}$ (i.e., the time the binary leaves it’s dissolving natal cluster). To avoid numerous intractable difficulties in the calculation, we instead adopt the conservative simplifying assumption of a constant disk density equal to the present-day density distribution.
@@
$\lambda_{\rm d,MW}=\frac{1}{v_{d}}\int^{R_{G}^{\prime}}_{R_{G}}\Gamma_{d}(R_{G})dR_{G}=\frac{22.7\,Ga\ln(\Lambda_{d})}{\sigma_{\rm MW}v_{d}}\int^{R_{G}^{\prime}}_{R_{G}}\rho_{MW}(R_{G})dR_{G},$
$v_{d}=\frac{R_{G}^{\prime}-R_{G}}{t^{\prime}},\ t^{\prime}=\tau_{\rm MW}-\tau_{\rm cl}-t$
where $v_{d}$ is the average “diffusion velocity” and $\tau_{\rm cl}$ is the total lifetime of the cluster.
@@
$\int^{R_{G}^{\prime}}_{R_{G}}\rho_{\rm MW}(R_{G})dR_{G}=$ $\ \sigma_{\rm disk}(M_{d},R_{G}^{\prime},c_{\rm ha},b_{\rm ha})-\sigma_{\rm disk}(M_{d},R_{G},c_{\rm ha},b_{\rm ha})$ (C3)
@@ C.1.2 Phase Mixing
After a binary escapes a cluster, if the binary is sufficiently small not to be destroyed by Galactic tides (i.e. if $a\lesssim 2\times 10^{5}$ AU), the binary evolves secularly under their influence (J. Heisler & S. Tremaine, 1986; C. Hamilton & R. R. Rafikov, 2019b, a; C. Hamilton, 2022; E. Grishin & H. B. Perets, 2022). C. Hamilton (2022) showed that the consequence of such evolution is a “phase mixing” (PM) in the appropriate phase-space region. This secular evolution is described by the Galactic-tide Hamiltonian $\mathscr{H}$, given by (C. Hamilton & R. R. Rafikov, 2019b, a)
[…]
@@ C.2 Cumulative Scatter
where $(a^{\prime},\epsilon^{\prime 2})$ are the updated binary properties during time step $\delta t$, $m_{b}=1M_{\odot}$ is the total mass of the binary, $m_{p}=0.5M_{\odot}$ is the average stellar mass, $v_{d}$ is the average radial diffusion velocity of the binary—see also Equation (C2)—and $(\xi_{a},\xi_{\epsilon})$ are independent Gaussian random numbers with mean 0 and variance 1.101010Note that in C. Hamilton & S. Modak (2024), $\epsilon$ is the square of the binary eccentricity, while this work employs $\epsilon$ as the eccentricity due to the symbolic degeneracy with the Euler’s number.
@@
where the expression has been simplified to the case of $1M_{\odot}$ binaries with a relative velocity dispersion between GMCs and disk stars of $\sim 30\rm{km\,s^{-1}}$. The probability that a binary evolving under CS survives interactions with GMCs in each time-step is
@@
$M_{\rm cl0}$ $50\ M_{\odot}$ $10^{6}\ M_{\odot}$ an initial mass containing 100 bodies to a mass encompassing ${\gg}99\%$ of the parameter space
@@
$R_{G}^{\prime}$ $7.9\ \rm kpc$ $8.1\ \rm kpc$ the subset of final orbital radii within 100 pc of the sun
@@ Appendix D The Galactic 3BBF Rate
Having assembled the creation, diffusion, and destruction portions of the final distribution, we are finally in a position to discuss subtleties of the total integration. The final equation, our “Galactic 3BBF Rate,” (G3R) is the product of Equations A10, A12, A16, B3, B4, and C2, built with the numerous intermittent expressions. Written in functional form, the G3R is
$\mathcal{G}_{\rm MW}d\Omega=\frac{\text{\makebox[0.0pt][l]{\ooalign{\hfil\hphantom{$V$}\hfil\cr\kern 0.80002pt--\hfil\cr}}}V_{*}}{\text{\makebox[0.0pt][l]{\ooalign{\hfil\hphantom{$V$}\hfil\cr\kern 0.80002pt--\hfil\cr}}}V_{\rm R_{G,s}}}f_{\rm CIMF}(M_{\rm cl0},t,R_{G})\mathcal{C}(M_{\rm cl0},\tilde{\tau},R_{G},r,k,\epsilon,a)P_{\rm no\,enc}(M_{\rm cl0},\tilde{\tau},R_{G},r,a)$ (D1)
$\qquad\qquad\times\mathcal{DHL}(t,\tau,R_{G},R_{G}^{\prime})P_{\rm d,MW}(a,t,\tau,R_{G},R_{G}^{\prime})d\Omega,$
@@
where
V
VRG,s
=δl2/(3RG,sh)=4.2×104
[…]
is the fraction of a $1\rm kpc$ thick disk annulus at $R_{\rm G,s}=8\rm kpc$ within a $\pm 100$ pc radius of the Sun, and the substitutions $r=r_{t}\,\tilde{r}$, $dr=r_{t}d\tilde{r}$, $\tau=\tau_{\rm cl}\,\tilde{\tau}$, and $d\tau=\tau_{\rm cl}d\tilde{\tau}$ are made for a more convenient integration experience.
@@
The integration bounds are detailed in Table 3. A number of additional constraints are applied to Equation (D1), setting it to 0 when triggered. These constraints include:
@@
    $N_{\rm cl}(\tilde{\tau})\leq 20$; the validity of a Plummer model is contingent to many bodies being present. We arbitrarily select $N=20$ as a break point for integration.
@@
    $a\left(1-\epsilon\right)<0.1\rm{AU}$; physical collisions and strong decay by tidal forces before and after 3BBF would prevent tight passages.
@@
    $\tau_{\rm MW}-\tau_{\rm cl}-t<0$; this only occurs if the migration timescale is larger than the time a binary has to migrate to $R_{G}^{\prime}$ from $R_{G}$ by the present day.
Incorporating all stated prior conditions with Equation (D1), the G3R is integrated over the entire parameter space using a custom written Monte Carlo integration scheme which divides the parameter space into $\sim 100$ sub-volumes and draws a total of $10^{9}$ samples, resolving the integral to high-precision. We additionally employ emcee (D. Foreman-Mackey et al., 2013) to uncover the underlying nine-dimensional distribution function accompanied by the same restrictions quoted above.